The elimination half-life of an antibiotic is 3 hours with an apparent volume of distribution equivalent to 20% of body weight. The usual therapeutic weight range for this antibiotic is between 5 and 15µg/ml. adverse toxicity for this drug is often observed at serum concentration greater than 20µg/ml. Calculate a dosage regimen (multiple IV doses) that will just maintain the serum drug concentration between 5 and 15µg/ml.
Answer #1

Let the weight of the patient is X kg.

Vd =20/100* X = 0.2 X L = 200 X ml

Target plasma concentration = 15 micrograms/ ml

Dose = 15 * 200= 3000 X micrograms = 3 X mg

So dose = 3mg/ kg

We have to read readminister the dose when the concentration falls to 5 micrograms/ml.

Half life = 3 hours

K = 0.693/3 = 0.231

A = Ao e^-kt

5 = 15 e^-0.231t

0.231 t = 1.1

t= 1.1/0.231 = 4.7 hours

So the dose should be administered every 4 and half hours.